OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0.

OpenStudy (anonymous):

A. slope = −1/4 and y-intercept = −9/4 B. slope = 1/4 and y-intercept = −9/4 C. slope = 1/4 and y-intercept = 9/4 D. slope = −1/4 and y-intercept = 9/4

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

B?

OpenStudy (triciaal):

i think C because x- 9/4 = 0 x = 9/4